OpenStudy (anonymous):

help!!!! triangle ABC is a right triangle.the sides measure AB=6 and BC=8.use the Pythagorean theorm to find CA. A.10 B.25 C.100 D.15

OpenStudy (diyadiya):

\[\huge AC^2 = AB^2 + BC^2\] This is the pythogorean theorem.

OpenStudy (anonymous):



OpenStudy (diyadiya):

Its basically states that squares of the legs of a right triangle add up to the square of the hypotenuse.

OpenStudy (anonymous):

is it c?

OpenStudy (diyadiya):

Can you show me how?

OpenStudy (diyadiya):

\[\Large AC^2 = AB^2 + BC^2 \]\[\Large AC^2 = (6)^2 + (8)^2 \] Remember \(\Large AC \ne AC^2\) :)

OpenStudy (anonymous):

thats what i did.so it has to equal more than 100?

OpenStudy (diyadiya):

\[\Large AC^2 = 100\] this is what you got ,But you have to find \(\Large AC \) Not \(\Large AC^2 \) How do you fins \(\Large AC \) Now?

OpenStudy (anonymous):

so its ten.because ten squared equals 100

OpenStudy (goformit100):

B.25 Ans.

OpenStudy (diyadiya):

Right!! \[\Large AC^2 = 100\]\[\Large \sqrt{AC^2} =\sqrt{100}\]\[\Large AC = 10\]

OpenStudy (diyadiya):

@goformit100 Please don't just give the final answer, If you think 25 is the answer, Please do explain. Thanks

OpenStudy (anonymous):

thank you!

OpenStudy (diyadiya):

You're Welcome :)

OpenStudy (anonymous):

can you help me with one more question?

OpenStudy (diyadiya):

I'll try ,Post it!

OpenStudy (anonymous):

okay hang on one second!